# Approximation Algorithms for Optimal Hopsets

### v1

@@ 5.1 Junction Tree Algorithm
To prove the theorem, we give an algorithm similar to a subroutine of the directed pairwise weighted spanner algorithm of [GKL23], where “weighted" refers to edge costs. Just as for hopsets, the directed pairwise weighted spanner problem does not have $n-1$ as a lower bound on the cost of the optimal solution. This allows their techniques to be useful in our setting.
In the pairwise weighted spanner subroutine of [GKL23], they define a variant of the junction tree (the “distance-preserving junction tree"). Junction trees are rooted trees that satisfy demands, and good junction trees satisfy many demands at low cost; that is, they have low “density." Junction trees have been used in several spanner approximation algorithms (e.g. [GKL23, CDKL20, GKL24]). In [GKL23], they give an algorithm that iteratively buys their version of low density junction trees until all demands are satisfied. Our algorithm will follow the same structure. The main technical work in this section is in showing that low-density hopbounded junction trees exist in our setting, and that we can use the subroutine of [GKL23] to find these hopbounded junction trees, even though their subroutine does not have any hop guarantees.
We first define a hopbounded variant of the junction tree, which we call an $(i,j)$-distance-preserving hopbounded junction tree. We parameterize by $i,j$, where $i+j\leq\beta$, to ensure that both “sides" of the rooted tree—the in-arborescence and the out-arborescence that make up the tree—are hopbounded by $i$ and $j$, respectively, so that all paths in the tree have at most $\beta$ hops.
@@
Going forward, we will refer to the $(i,j)$-distance-preserving hopbounded junction tree as simply an “$(i,j)$-junction tree." Our algorithm will find and remove a low-density hopbounded junction tree from $G_{M}$, add its edges to the current solution, and repeat, until all demand pairs are settled. We will give an $O(n^{\epsilon})$-approximation for finding these low-density junction trees. The algorithm will return a subgraph with total cost of $\widetilde{O}(\beta^{2}n^{\epsilon}\cdot{{{\sf OPT}}}^{2})$.
@@ 5.1.2 Layered Graph Reduction
For any $(i,j)$-junction tree $T$ in $G_{M}$, there exists an $(i,j)$-junction tree $T_{L}$ in $G_{L}$ with the same density.
@@
Finally, we add “dummy paths" to $T_{L}$ to ensure that the corresponding demands in $\mathcal{D}_{L}$ are settled. These dummy paths are need to handle demand-settling paths in $T$ that have fewer that $\beta$ hops. Let $A^{in}_{L}$ and $A^{out}_{L}$ denote the in- and out-arborescences of $T_{L}$, respectively. For each vertex $v_{x}\in A^{in}_{L}$, we add edges $\{(v_{k},v_{k+1}):0\leq k<x\}$ (and corresponding nodes) to $T_{L}$ (if they don’t already exist in $T_{L}$). Similarly for each vertex $v_{x}\in A^{out}_{L}$, we add edges $\{(v_{k},v_{k+1}):x\leq k<\beta\}$ (and corresponding nodes) to $T_{L}$ (if they don’t already exist in $T_{L}$). With these dummy paths, we have that if demand $(s,t)$ is settled by a path in $T$ with fewer than $\beta$ hops, then the corresponding demand $(s_{0},t_{\beta})$ is also settled in $T_{L}$.
@@
The number of demands satisfied in both trees is also the same, which we show by mapping each demand $(s,t)\in\mathcal{D}$ settled by $T$ to a unique demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$ (and vice versa). We now show that if $(s,t)$ is settled by $T$, then $(s_{0},t_{\beta})$ is settled by $T_{L}$. Let $P=\{s,a,b,\dots,t\}$ be the path in $T$ that settles $(s,t)$. Then, path $P_{L}=\{s_{0},\dots,s_{k},a_{k+1},b_{k+2},\dots,t_{m},\dots,t_{\beta}\}$ is in $T_{L}$, for some $k,m$. Paths $P$ and $P_{L}$ have the same length—any edge $(u,v)\in P$ has the same length as its corresponding edge $(u_{k},v_{k+1})$ (for some $k$) in $P_{L}$, and all other edges (i.e., edges from the dummy subpaths $(s_{0},\dots,s_{k})$ and $(t_{m},\dots,t_{\beta})$, if they exist) have length 00. We therefore have that $d_{T_{L}}(s_{0},t_{\beta})=d_{T}(s,t)\leq Dist(s,t)=Dist_{L}(s_{0},t_{\beta})$. Also, path $P_{L}$ has exactly $\beta$ hops. Thus, demand $(s_{0},t_{\beta})$ is settled by $T_{L}$. It is also not difficult to see that each demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$ can be mapped to a unique demand $(s,t)\in\mathcal{D}$ settled by $T$, using similar arguments. Trees $T$ and $T_{L}$ therefore have the same cost and settle the same number of demands, and so have the same density. ∎
@@
For any $(i,j)$-junction tree $T_{L}$ in $G_{L}$, there exists an $(i,j)$-junction tree $T$ in $G_{M}$ with the same density.
@@
Both trees also settle the same number of demands. Each demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$ can be mapped to a unique demand $(s,t)\in\mathcal{D}$ settled by $T$. Let $P_{L}=\{s_{0},\dots,s_{k},a_{k+1},b_{k+2},\dots,t_{m},\dots,t_{\beta}\}$ be the path that settles $(s_{0},t_{\beta})$ in $T_{L}$. Then, the path $P=\{s,a,b,\dots,t\}$ is in $T$. Paths $P$ and $P_{L}$ have the same length—any edge of the form $(u_{k},v_{k+1})\in P_{L}$ (for some $k$), where $u\neq v$, has the same length as its corresponding edge $(u,v)\in P$. All other edges in $P_{L}$ (i.e., edges from the dummy subpaths if they exist) have length 00. We therefore have that $d_{T}(s,t)=d_{T_{L}}(s_{0},t_{\beta})\leq Dist_{L}(s_{0},t_{\beta})=Dist(s,t)$. Path $P$ also has at most $\beta$ hops, so $(s,t)$ is settled by $T$. It is also not difficult to see that each demand $(s,t)\in\mathcal{D}$ settled by $T$ can be mapped to a unique demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$, using similar arguments. We’ve shown that $T_{L}$ and $T$ have the same cost and settle the same number of demands, and so they have the same density. ∎
@@ 5.1.3 Junction Tree-Finding Subroutine
By iteratively buying these low-density $(i,j)$-junction trees, we get an $O(\beta n^{\epsilon}\cdot{{{\sf OPT}}}^{2})$-approximation for Generalized $\beta$-Hopset.
@@ 5.2 Star Sampling with Randomized LP Rounding Algorithm
The pair of algorithms we give closely follow the $\widetilde{O}(n^{2/3})$- and $\widetilde{O}(\sqrt{n})$-approximations for the unweighted $k$-spanner problem, given by [DK11] and [BBM+11]. The $k$-spanner algorithm is a trade-off between two algorithms: an arborescence sampling algorithm for settling a class of edges (or demands) that they call “thick," and a randomized LP rounding algorithm for settling “thin" edges. For hopsets we will settle these thick demands by sampling directed stars instead of arborescences, and for thin demands we will use a similar LP rounding approach. Although our hopset algorithms are similar to the $k$-spanner algorithms, we get a different approximation ($\widetilde{O}(n\big{/}\sqrt{{\sf OPT}})$ for hopsets versus $\widetilde{O}(\sqrt{n})$ for spanners). This is because [DK11, BBM+11] take advantage of the fact that for spanners, ${{\sf OPT}}\geq n-1$. This is not the case for hopsets so we get a different approximation out of the algorithms, in terms of OPT.
We note that our $O(n\ln{n}\,\big{/}\sqrt{{\sf OPT}})$-approximation is achieved by trading off the two aforementioned algorithms (star-sampling and randomized-rounding). To later achieve the optimal trade-off with other algorithms, one should a priori treat each of these two algorithms as separate, with their own individual approximation ratios. It is however equivalent to trade these two algorithms off first and treat them as one combined algorithm, which we do going forward. This is because these are our only algorithms that will depend on the “local neighborhood size" parameter.
@@ 5.2.1 Star-Sampling Algorithm for Thick Demands
We now show that $E^{\prime}$ has the desired cost in expectation. While [DK11] proves this for spanners, it’s easy to see that a near identical argument also holds for hopsets in $G_{M}$. We restate the proof for completeness.
@@ 5.2.2 Randomized LP Rounding Algorithm for Thin Demands
We now give the algorithm for finding a set $E^{\prime\prime}$ to settle thin demands. [BBM+11] introduces the notion “anti-spanners," which is crucial for the algorithm and analysis for settling thin demands. In particular, they formulate an anti-spanner covering LP that captures the problem of settling all thin demands. They then solve the LP (with high probability) by constructing a separation oracle that utilizes randomized rounding. We will also use randomized LP rounding, though instead of rounding the solution to an “anti-hopset" covering LP, we will round based on LP(Hopset). Our LP is stronger than the “anti-hopset" covering LP, since our LP is for fractional cuts against valid paths, while the anti-hopset covering LP is only for integer cuts.
@@
We use this modified version of LP(Hopset) for the randomized rounding algorithm. Given a fractional solution $\bm{\mathrm{x}}^{*}$ to LP(Hopset), our algorithm will return an edge set $E^{\prime\prime}$ that, with high probability, will cost at most $2{{\sf OPT}}\cdot 2(n/b)\ln{n}$ and satisfy all thin demands (see Algorithm 3). We say that the algorithm fails if $c(E^{\prime\prime})>2{{\sf OPT}}\cdot 2(n/b)\ln{n}$ or if $E^{\prime\prime}$ doesn’t satisfy all thin demands. The algorithm will fail with low probability.
@@
To show that with high probability, Algorithm 3 does not fail, we start by defining “anti-hopsets," the analogous of anti-spanners in the hopset setting. For a given demand $(s,t)$, an anti-hopset is a set of edges such that removing them from $G_{M}$ results in no valid paths from $s$ to $t$ in what remains.
@@
The probability that the algorithm fails is exponentially small in $n$.
@@
If ${{\sf OPT}}=\widetilde{O}(n^{2})$, then we can achieve an $\widetilde{O}(1)$-approximation by just returning $E_{M}\setminus E^{\prime\prime}$. Thus, we can assume without loss of generality that ${{\sf OPT}}=\widetilde{o}(n^{2})$. We therefore have that the probability that $E^{\prime\prime}$ doesn’t cover all minimal anti-hopsets for a thin demand $(s,t)$ is exponentially small. ∎
@@ 5.3 Trade-Offs with Existential Bounds
We first state a folklore existential bound that we will use throughout. For exact hopsets in both directed and undirected graphs, the following bound is the best-known (and is known to be tight [BH23]). Exact hopsets satisfy any arbitrary stretch, so we can trade off the hopset produced by the folklore existential bound with other algorithms in any regime. For more restricted regimes—specifically for limited stretch, hopbound, and undirected graphs—we will trade off with stronger existential bounds to get improved approximations.

### v2

@@ 5.1 Junction Tree Algorithm
To prove the theorem, we give an algorithm similar to a subroutine of the directed pairwise weighted spanner algorithm of [GKL23], where “weighted” refers to edge costs. Just as for hopsets, the directed pairwise weighted spanner problem does not have $n-1$ as a lower bound on the cost of the optimal solution. This allows their techniques to be useful in our setting.
In the pairwise weighted spanner subroutine of [GKL23], they define a variant of the junction tree (the “distance-preserving junction tree”). Junction trees are rooted trees that satisfy demands, and good junction trees satisfy many demands at low cost; that is, they have low “density.” Junction trees have been used in several spanner approximation algorithms (e.g. [GKL23, CDKL20, GKL24b]). In [GKL23], they give an algorithm that iteratively buys their version of low density junction trees until all demands are satisfied. Our algorithm will follow the same structure. The main technical work in this section is in showing that low-density hopbounded junction trees exist in our setting, and that we can use the subroutine of [GKL23] to find these hopbounded junction trees, even though their subroutine does not have any hop guarantees.
We note that the junction tree framework developed by [GKL24b] for multicriteria spanners also works for our setting. Their framework implies an $\widetilde{O}(|\mathcal{D}|^{\epsilon}\cdot\beta)$-approximation for finding the hopbounded minimum-density junction tree, and thus a $\widetilde{O}(\beta^{2}\cdot|\mathcal{D}|^{\epsilon}\cdot{{\sf OPT}})$-approximation for directed Generalized $\beta$-Hopset. By tailoring our framework to hopsets, we achieve an $O(n^{\epsilon})$-approximation for finding the hopbounded min-density junction tree (note the lack of dependence on $\beta$), implying our main result of this section (Theorem 5.1), a $\widetilde{O}(\beta\cdot n^{\epsilon}\cdot{{\sf OPT}})$-approximation overall. By using our hopbounded junction tree framework, we lose a factor $\beta$ in our overall approximation compared to what is implied by [GKL24b].
We first define a hopbounded variant of the junction tree, which we call an $(i,j)$-distance-preserving hopbounded junction tree. We parameterize by $i,j$, where $i+j\leq\beta$, to ensure that both “sides” of the rooted tree—the in-arborescence and the out-arborescence that make up the tree—are hopbounded by $i$ and $j$, respectively, so that all paths in the tree have at most $\beta$ hops.
@@
Going forward, we will refer to the $(i,j)$-distance-preserving hopbounded junction tree as simply an “$(i,j)$-junction tree.” Our algorithm will find and remove a low-density hopbounded junction tree from $G_{M}$, add its edges to the current solution, and repeat, until all demand pairs are settled. We will give an $O(n^{\epsilon})$-approximation for finding these low-density junction trees. The algorithm will return a subgraph with total cost of $\widetilde{O}(\beta^{2}n^{\epsilon}\cdot{{{\sf OPT}}}^{2})$.
@@ 5.1.2 Layered Graph Reduction
For any $(i,j)$-junction tree in $G_{M}$, there exists an $(i,j)$-junction tree in $G_{L}$ with the same density.
@@
Finally, we add “dummy paths” to $T_{L}$ to ensure that the corresponding demands in $\mathcal{D}_{L}$ are settled. These dummy paths will handle demand-settling paths in $T$ that have fewer that $\beta$ hops. Let $A^{in}_{L}$ and $A^{out}_{L}$ denote the in- and out-arborescences of $T_{L}$, respectively. For each vertex $v_{x}\in A^{in}_{L}$, we add edges $\{(v_{k},v_{k+1}):0\leq k<x\}$ (and corresponding nodes) to $T_{L}$ (if they don’t already exist in $T_{L}$). Similarly for each vertex $v_{x}\in A^{out}_{L}$, we add edges $\{(v_{k},v_{k+1}):x\leq k<\beta\}$ (and corresponding nodes) to $T_{L}$ (if they don’t already exist in $T_{L}$). With these dummy paths, we ensure that if demand $(s,t)$ is settled by a path in $T$ with fewer than $\beta$ hops, then the corresponding demand $(s_{0},t_{\beta})$ is also settled in $T_{L}$.
@@
The number of demands satisfied in both trees is also the same, which we show by mapping each demand $(s,t)\in\mathcal{D}$ settled by $T$ to a unique demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$ (and vice versa). We now show that if $(s,t)$ is settled by $T$, then $(s_{0},t_{\beta})$ is settled by $T_{L}$. Let $P=(s,a,b,c\dots,t)$ be the path in $T$ that settles $(s,t)$. Then, path $P_{L}=(s_{0},\dots,s_{k},a_{k+1},b_{k+2},c_{k+3},\dots,t_{m},\dots,t_{\beta})$ is in $T_{L}$, for some $k,m$. Note that subpaths $(s_{0},\dots,s_{k})$ and $(t_{m},\dots,t_{\beta})$ are dummy subpaths. Paths $P$ and $P_{L}$ have the same length—any edge $(u,v)\in P$ has the same length as its corresponding edge $(u_{k},v_{k+1})$ (for some $k$) in $P_{L}$, and all other edges (i.e., edges from the dummy subpaths $(s_{0},\dots,s_{k})$ and $(t_{m},\dots,t_{\beta})$, if they exist) have length 00. We therefore have that $d_{T_{L}}(s_{0},t_{\beta})=d_{T}(s,t)\leq Dist(s,t)=Dist_{L}(s_{0},t_{\beta})$. Also, path $P_{L}$ has exactly $\beta$ hops. Thus, demand $(s_{0},t_{\beta})$ is settled by $T_{L}$. It is also not difficult to see that each demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$ can be mapped to a unique demand $(s,t)\in\mathcal{D}$ settled by $T$, using similar arguments. Trees $T$ and $T_{L}$ therefore have the same cost and settle the same number of demands, and so have the same density. ∎
@@
For any $(i,j)$-junction tree in $G_{L}$, there exists an $(i,j)$-junction tree in $G_{M}$ with the same density.
@@
Both trees also settle the same number of demands. Each demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$ can be mapped to a unique demand $(s,t)\in\mathcal{D}$ settled by $T$. Let $P_{L}=(s_{0},\dots,s_{k},a_{k+1},b_{k+2},c_{k+3},\dots,t_{m},\dots,t_{\beta})$ be the path that settles $(s_{0},t_{\beta})$ in $T_{L}$. Then, the path $P=(s,a,b,c,\dots,t)$ is in $T$. Paths $P$ and $P_{L}$ have the same length—any edge of the form $(u_{k},v_{k+1})\in P_{L}$ (for some $k$), where $u\neq v$, has the same length as its corresponding edge $(u,v)\in P$. All other edges in $P_{L}$ (i.e., edges from the dummy subpaths if they exist) have length 00. We therefore have that $d_{T}(s,t)=d_{T_{L}}(s_{0},t_{\beta})\leq Dist_{L}(s_{0},t_{\beta})=Dist(s,t)$. Path $P$ also has at most $\beta$ hops, so $(s,t)$ is settled by $T$. It is also not difficult to see that each demand $(s,t)\in\mathcal{D}$ settled by $T$ can be mapped to a unique demand $(s_{0},t_{\beta})\in\mathcal{D}_{L}$ settled by $T_{L}$, using similar arguments. We’ve shown that $T_{L}$ and $T$ have the same cost and settle the same number of demands, and so they have the same density. ∎
@@ 5.1.3 Junction Tree-Finding Subroutine
By iteratively buying these low-density $(i,j)$-junction trees, we get an $O(\beta n^{\epsilon}\cdot{{{\sf OPT}}})$-approximation for Generalized $\beta$-Hopset.
@@ 5.2 Star Sampling with Randomized LP Rounding Algorithm
The pair of algorithms we give closely follow the $\widetilde{O}(n^{2/3})$- and $\widetilde{O}(\sqrt{n})$-approximations for the unweighted $k$-spanner problem, given by [DK11] and [BBM+11], respectively. The $k$-spanner algorithm is a trade-off between two algorithms: an arborescence sampling algorithm for settling a class of edges (or demands) that they call “thick,” and a randomized LP rounding algorithm for settling “thin” edges. For hopsets we will settle these thick demands by sampling directed stars instead of arborescences, and for thin demands we will use a similar LP rounding approach. Although our hopset algorithms are similar to the $k$-spanner algorithms, we get a different approximation ($\widetilde{O}(n\big{/}\sqrt{{\sf OPT}})$ for hopsets versus $\widetilde{O}(\sqrt{n})$ for spanners). This is because [DK11, BBM+11] take advantage of the fact that for spanners, ${{\sf OPT}}\geq n-1$. This is not the case for hopsets so we get a different approximation out of the algorithms, in terms of OPT. This approach is also similar to that of [CDKL20] for pairwise distance preservers, where again, $\Omega(n)$ is not a lower bound for OPT.
We note that our $O(n\ln{n}\,\big{/}\sqrt{{\sf OPT}})$-approximation is achieved by trading off the two aforementioned algorithms (star-sampling and randomized-rounding). To later achieve the optimal trade-off with other algorithms, one should a priori treat each of these two algorithms as separate, with their own individual approximation ratios. It is however equivalent to trade these two algorithms off first and treat them as one combined algorithm, which we do going forward. This is because these are our only algorithms that will depend on the “local neighborhood size” parameter.
@@ 5.2.1 Star-Sampling Algorithm for Thick Demands
We now show that $E^{\prime}$ has the desired cost in expectation. While [DK11] proves this for spanners, it is easy to see that a near identical argument also holds for hopsets in $G_{M}$. We restate the proof for completeness.
@@ 5.2.2 Randomized LP Rounding Algorithm for Thin Demands
We now give the algorithm for finding a set $E^{\prime\prime}$ to settle thin demands. [BBM+11] introduces the notion “anti-spanners,” which is crucial for the algorithm and analysis for settling thin demands. In particular, they formulate an anti-spanner covering LP that captures the problem of settling all thin demands. They then solve the LP (with high probability) by constructing a separation oracle that utilizes randomized rounding. We will also use randomized LP rounding, though instead of rounding the solution to an “anti-hopset” covering LP, we will round based on LP(Hopset). Our LP is stronger than the “anti-hopset” covering LP, since our LP is for fractional cuts against valid paths, while the anti-hopset covering LP is only for integer cuts.
@@
We use this modified version of LP(Hopset) for the randomized rounding algorithm. Given a fractional solution $\bm{\mathrm{x}}^{*}$ to LP(Hopset), our algorithm will return an edge set $E^{\prime\prime}$ that, with high probability, will cost at most $2{{\sf OPT}}\cdot 2(n/b)\ln{n}$ and satisfy all thin demands (see Algorithm 3). We say that the algorithm fails if $c(E^{\prime\prime})>2{{\sf OPT}}\cdot 2(n/b)\ln{n}$ or if $E^{\prime\prime}$ does not satisfy all thin demands. The algorithm will fail with low probability.
@@
To show that with high probability, Algorithm 3 does not fail, we start by defining “anti-hopsets,” the analogous of anti-spanners in the hopset setting. For a given demand $(s,t)$, an anti-hopset is a set of edges such that removing them from $G_{M}$ results in no valid paths from $s$ to $t$ in what remains.
@@
The probability that Algorithm 3 fails is exponentially small in $n$.
@@
If ${{\sf OPT}}=\widetilde{\Theta}(n^{2})$, then we can achieve an $\widetilde{O}(1)$-approximation by just returning $E_{M}\setminus E^{\prime\prime}$. Thus, we can assume without loss of generality that ${{\sf OPT}}=\widetilde{o}(n^{2})$. Given this, we have that the probability that $E^{\prime\prime}$ does not cover all minimal anti-hopsets for a thin demand $(s,t)$ is exponentially small. ∎
@@ 5.3 Trade-Offs with Existential Bounds
We first state a folklore existential bound that we will use throughout. For exact hopsets in both directed and undirected graphs, the following bound is the best-known (and is known to be tight [BH23]). Exact hopsets satisfy any distance bound function, so we can trade off the hopset produced by the folklore existential bound with other algorithms in any regime. For more restricted regimes—specifically for limited stretch, hopbound, and undirected graphs—we will trade off with stronger existential bounds to get improved approximations.